Legess Legacy Company produces machine. The company incurs the following monthly costs to produces the machine.

Fixed costs = RM 350000

Variable costs = RM 4000 per unit

The company produces 200 units per month and the machine sell for RM 8000 per unit.

 a. Determine:

     i. the total cost, total revenue and total profit for the company.

     ii. break even volume for the company.

   b. Determine the new break even volume if the company:

  1. reduce the raw material cost by 10%.
  2. increase the price of the machine by 5%.
1
Expert's answer
2021-06-27T04:03:56-0400

Solution.

a.

We have Fixed costs

FC = 350000

Variable costs

VC=4000 per unit

Quantity Q=200 units

Price P=8000 per unit.

Therefore,

Total cost is TC=FC+VC×Q= 350000+4000×200=1150000.

Total revenue is TR=P×Q=8000×200=1600000.

Total profit is TP=TR-TC=1600000-1150000=450000.

Break even volume for the company is "\frac{TC}{P-VC}=\frac{350000}{8000-4000}=87.5"

So, break even volume for the company is equal to 88 machines.

b.

If the company reduces the raw material cost by 10%, then

4000-(4000:100×10)=4000-400=3600 per unit.

Break even volume for the company in this case will be

"\frac{350000}{8000-3600}=79.5"

So, break even volume for the company is equal to 80 machines.

If the company increases the price of the machine by 5%, then

8000+(8000:100×5)=8000+400=8400 per machine.

Break even volume for the company in this case will be

"\frac{350000}{8400-4000}=79.5"

So, break even volume for the company is equal to 80 machines.